# Agglomerative Clustering of Simulation Output Distributions Using Regularized Wasserstein Distance

### v1

@@ Abstract
We investigate the use of clustering methods on data produced by a stochastic simulator, with applications in anomaly detection, pre-optimization, and online monitoring. We introduce an agglomerative clustering algorithm that clusters multivariate empirical distributions using the regularized Wasserstein distance and apply the proposed methodology on a call-center model.
@@ 1 Introduction
A simulation model is a computational or mathematical representation of a real-world system designed to study its behavior under various scenarios. Simulation models are used extensively in fields such as engineering, economics, healthcare, and environmental science to predict and analyze the outcomes of different scenarios without the need to experiment in the real world, which can be costly, time-consuming, or impractical. Outputs of a simulation model typically correspond to key performance indicators (KPIs) of interest to the decision maker, e.g., profit, throughput, or service level. For stochastic simulation models, simulating a given scenario generates outputs that vary from replication to replication, thus each scenario has an associated probability distribution describing the stochastic behavior of its outputs. When conducting simulation experiments, the user controls which scenarios are simulated and how many replications are run. Simulation experiments can be easily designed to generate data that satisfies the standard assumptions of being independent and identically distributed, in contrast to most data obtained from the real world. Common tools for analyzing simulation output data include summary statistics (e.g., sample means, variance, and covariances) and visualization tools (e.g., histograms and boxplots). For problems with multiple KPIs, the multivariate empirical distribution produced by the data contains valuable information about system performance, but can be difficult to analyze and plot.
To reveal important patterns and relationships that cannot be detected by conventional data analysis methods, we propose clustering the empirical distributions of simulated scenarios.
Clustering is an unsupervised learning approach that can help discover important patterns and relationships in complex datasets. In the context of simulation output data, clustering can identify scenarios with similar KPIs, or more precisely, similar output distributions. Moreover, clustering facilitates comparative analysis—by understanding the characteristics of each cluster, a decision maker can draw meaningful comparisons and make informed decisions.
We consider three important applications of clustering for enhancing simulation output analysis: anomaly detection, which involves identifying and investigating outliers in simulation outputs; pre-optimization, which involves formulating simulation-optimization problems and identifying promising initial solutions; and online monitoring, which involves tracking the system state over time and using classification methods to detect potentially undesirable system behavior and trigger appropriate actions or alerts.
@@
We propose an agglomerative clustering method that uses the complete-linkage criterion for forming clusters. We choose agglomerative clustering for its flexibility, as it does not require specifying a predetermined number of clusters, and opt for complete linkage because it maintains compact and well-separated clusters. For the algorithm’s measure of dissimilarity between distributions, we choose the regularized Wasserstein distance. The (unregularized) Wasserstein distance quantifies the distance between two discrete probability distributions by the minimum amount of “work”—defined as the product of the probability mass that needs to be moved and the distance it needs to be transported—required to transform one distribution into the other (Villani, , 2009). The Wasserstein distance is chosen over other metrics like Kullback-Leibler divergence or Jensen-Shannon divergence due to its ability to handle distributions with non-overlapping supports, as arises when working with continuous-valued empirical distributions. Another advantage of the Wasserstein distance is the notion of a barycenter, which acts like an average among distributions. This property is particularly useful for post-hoc analysis, when desiring to summarize each cluster with a representative distribution (Peyré & Cuturi, , 2019; Arjovsky et al., , 2017; Cuturi, , 2013). However, computing the Wasserstein distance between two discrete distributions entails solving a linear program (Villani, , 2009), which could become computationally intensive when working with large datasets. The regularized Wasserstein distance, also known as the Sinkhorn distance, adds an entropic regularization term that promotes smoother transport plans, is easier to compute, and more stable, making it well suited for data-intensive applications (Cuturi, , 2013). Our algorithm determines the optimal number of clusters based on the silhouette index (Shahapure & Nicholas, , 2020), a centroid-free metric that evaluates clustering quality in terms of intra-cluster and inter-cluster distances.
@@
Clustering has been used for many purposes in many fields, such as analyzing gene expression in bioinformatics (Eisen et al., , 1998), market segmentation in economics (Wedel & Kamakura, , 2000), and textures and shapes in image processing (Pappas & Jayant, , 1989). Hierarchical clustering techniques, specifically agglomerative clustering, build nested clusterings by iteratively merging the two most similar clusters (Murtagh & Contreras, , 2012) and produce a dendrogram that depicts the nested clusterings at different levels (Jain et al., , 1999). Adaptations like single, complete, and average linkage offer great flexibility in how agglomerative clustering algorithms merge clusters (Jain, , 2010). More recently, clustering has been employed in the study of stochastic simulation models, including for simulation optimization (Li et al., , 2024; Zhang & Peng, , 2024; Peng et al., , 2018) and reducing model-form uncertainty (Abdallah et al., , 2020). To the best of our knowledge, we are the first to propose agglomerative clustering of simulation output distributions and investigate important use cases.
@@
The Wasserstein distance is used extensively in machine learning applications, such as in computer vision and pattern recognition to robustly compare visual feature distributions (Rubner et al., , 2000). For large-scale problems, variations like the entropy-regularized Wasserstein distance improve computational efficiency. For example, Benamou et al., (2015) employ iterative Bregman projections to efficiently solve regularized transportation problems, thereby improving scalability and accuracy. Our clustering framework uses the algorithm of Benamou et al., (2015) to calculate the regularized Wasserstein distance and regularized Wasserstein barycenters. For a more detailed overview of computational techniques for optimal transport and its practical applications, we direct the reader to Peyré & Cuturi, (2019).
The Wasserstein distance has previously appeared in $k$-means clustering, e.g., for clustering market regimes (Horvath et al., , 2021) and financial data (Riess et al., , 2023). Henderson et al., (2015) extend the classical $k$-means algorithm to the clustering of one-dimensional, continuous-valued empirical distributions. However, their clustering algorithm, EP-MEANS, becomes computationally inefficient as the sample size increases and is difficult to extend to multivariate empirical distributions. Zhuang et al., (2022) generalize the distance-based formulation of $k$-means to the Wasserstein space but identify several shortcomings, including the irregularity and non-robustness of barycenter-based Wasserstein $k$-means. To address scalability issues when clustering discrete distributions, Ye et al., (2017) approximate discrete Wasserstein barycenters for large clusters using a modified Bregman alternating direction method of multipliers (ADMM) approach. Our hierarchical clustering algorithm avoids many of these regularity and scalability issues. Chakraborty et al., (2020) introduce a hierarchical clustering algorithm that utilizes optimal transport-based distance measures, though the instances being clustered are not themselves distributions, as in our setting.
@@ 2 Use Cases
We first discuss the utility and versatility of clustering simulation output distributions through several use cases.
@@
In the context of simulation experiments, anomalies can be categorized as artificial or systemic. An artificial anomaly is typically associated with logic or coding errors within the simulation model, whereas a systemic anomaly is related to inherent features of the system. When using hierarchical clustering algorithms, anomalous output distributions can be identified by examining the dendrogram, the distances between clusters, or the cluster sizes (Loureiro et al., , 2004). After identifying an anomalous output distribution, one would first scrutinize the simulation code to determine if it is an artificial anomaly. If the anomaly is not artificial, then one might investigate further by, for example, examining the marginal distributions, correlation matrices, and corresponding input variables. Hierarchical clustering algorithms such as ours are expected to be more stable for identifying outliers than non-hierarchical clustering methods whose clusterings strongly depend on the initialization of the clusters.
@@
In many practical situations, there are tradeoffs between multiple KPIs, and the decision maker may be unable to articulate a priori what constitutes “good” versus “bad” system performance. By clustering output distributions and obtaining the barycenters, the decision maker can be presented with a more manageable number of distributions to compare. The decision maker could then conduct a series of A/B comparisons, wherein pairs of barycenters are compared and some are eliminated based on unformalized notions of preferred performance until a small number of clusters (or scenarios) remain. The clustering analysis can also help the decision maker specify which metrics should be modeled as objectives in a subsequent simulation-optimization problem and which should be treated as constraints. Achievable thresholds for the constraints can be set based on the observed performance outcomes of the simulated scenarios. Additionally, by examining the inputs associated with scenarios in a promising cluster, the decision maker can identify promising regions of the input space from which to initiate an optimization search, potentially leading to more rapid progress toward the optimal solution.
@@
This application concerns how the output of a simulation model is influenced by state variables, namely those that evolve over time and can be observed, but not directly controlled, by the decision maker. We envisage an online monitoring framework in which clustering is performed offline and state variables are later tracked in real time with classification algorithms being utilized to help the decision maker anticipate changes in system performance. This approach involves a preliminary simulation experiment in which the scenarios correspond to different initial states, followed by the clustering of the generated outputs. When monitoring the system’s state online, classification algorithms can be used to predict the cluster to which an observed state’s output distribution may belong. Conversely, if the classification algorithm struggles to assign a state to a single cluster, such as a tie when using $k$-nearest neighbors, it would suggest that system performance may change soon, potentially prompting intervention.
@@ 3 Clustering Simulation Output Distributions
Suppose there are $N$ scenarios under consideration, and for each Scenario $i$, $i=1,2,\ldots,N$, we obtain $n_{i}$ independent simulation replications. Let $\mathbf{y}_{il}\in\mathbb{R}^{d}$ denote the vector output of the $l$th simulation replication at Scenario $i$ and let $\mu_{i}:=n_{i}^{-1}\sum_{l=1}^{n_{i}}\delta_{y_{il}}$ denote the corresponding empirical distribution, i.e., a discrete probability distribution with support $\mathcal{Y}_{i}=\{\mathbf{y}_{i1},\ldots,\mathbf{y}_{in_{i}}\}$, where we ignore any duplicate values in the definition of $\mathcal{Y}_{i}$, and $\delta_{y_{il}}$ is the Dirac delta function at $y_{il}$. Since we are interested in clustering $\{\mu_{1},\mu_{2},\ldots,\mu_{N}\}$, and there is no specific ordering among distributions, we henceforth drop the subscript $i$ and denote the probability mass vector, support, and cardinality of the support of an empirical distribution $\mu$ as $\mathbf{p}_{\mu}$, $\mathcal{Y}_{\mu}$, and $M_{\mu}=|\mathcal{Y}_{\mu}|$, respectively.
@@ 3.1 Wasserstein Distance
Let $\Delta_{M}:=\left\{\mathbf{p}\in\mathbb{R}_{+}^{{M}}\colon\sum_{l=1}^{{M}}p_{l}=1\right\}$ denote the set of all possible probability mass vectors on a support of size $M$. For two empirical distributions $\mu$ and $\mu^{\prime}$ having probability mass vectors $\mathbf{p}_{\mu}\in\Delta_{M_{\mu}}$ and $\mathbf{p}_{\mu^{\prime}}\in\Delta_{M_{\mu^{\prime}}}$, respectively, the polytope of couplings is defined as
@@
The regularized Wasserstein distance, on the other hand, can be solved in near-linear time (Cuturi, , 2013) and is defined as
$W_{\lambda}(\mu,\mu^{\prime}):=\underset{\boldsymbol{\gamma}_{\lambda}\in\Pi(p_{\mu},p_{\mu^{\prime}})}{\operatorname{min}}\langle\mathbf{D},\boldsymbol{\gamma}_{\lambda}\rangle-\lambda E(\boldsymbol{\gamma}_{\lambda}),$ (2)
where $\lambda$ is a regularization parameter, and $E(\boldsymbol{\gamma}_{\lambda})$ is the entropy of the transportation plan matrix $\boldsymbol{\gamma}_{\lambda}$, defined as $E(\boldsymbol{\gamma}):=-\sum_{l=1}^{M_{\mu}}\sum_{l^{\prime}=1}^{M_{\mu^{\prime}}}\boldsymbol{\gamma}_{ll^{\prime}}\log\boldsymbol{\gamma}_{ll^{\prime}}$, where we set $\boldsymbol{\gamma}_{ll^{\prime}}\log\boldsymbol{\gamma}_{ll^{\prime}}=0$ if $\boldsymbol{\gamma}_{ll^{\prime}}=0$. As $\lambda$ approaches 00, the optimal transportation plan matrix for (2), $\boldsymbol{\gamma}_{\lambda}^{*}$, becomes more sparse and approaches $\boldsymbol{\gamma}^{*}$ (Benamou et al., , 2015). The entropic regularization term incentivizes $\boldsymbol{\gamma}_{\lambda}^{*}$ to be more diffuse than $\boldsymbol{\gamma}^{*}$, and this induced non-sparsity helps to stabilize the computation of $\boldsymbol{\gamma}_{\lambda}^{*}$ because (2) is a strongly convex program with a unique solution. An advantage of the regularized Wasserstein distance is that $\boldsymbol{\gamma}^{*}_{\lambda}$ can be calculated through an efficient iterative procedure involving matrix multiplications, as described in Algorithm 1.
Algorithm 1 Regularized Wasserstein distance (Benamou et al., , 2015)
@@
2:Construct a matrix $\mathbf{Q}\in\mathbb{R}^{M_{\mu}\times M_{\mu^{\prime}}}$ having entries $Q_{ll^{\prime}}=e^{-\frac{D_{ll^{\prime}}}{\lambda}}$ for $l=1,2,\ldots,M_{\mu}$ and $l^{\prime}=1,2,\ldots,M_{\mu^{\prime}}$.
@@
The stopping criteria in Algorithm 1 helps to control the computational cost and could involve setting a maximum number of iterations or stopping when the percentage change in the regularized Wasserstein distance is less than some threshold. The regularized Wasserstein distance plays a central role in our proposed clustering algorithm.

### v2

# An Agglomerative Clustering Algorithm for Simulation Output Distributions Using Regularized Wasserstein Distance

@@ Abstract
Using statistical learning methods to analyze stochastic simulation outputs can significantly enhance decision-making by uncovering relationships between different simulated systems and between a system’s inputs and outputs. We focus on clustering multivariate empirical distributions of simulation outputs to identify patterns and trade-offs among performance measures. We present a novel agglomerative clustering algorithm that utilizes the regularized Wasserstein distance to cluster these multivariate empirical distributions. This framework has several important use cases, including anomaly detection, pre-optimization, and online monitoring. In numerical experiments involving a call-center model, we demonstrate how this methodology can identify staffing plans that yield similar performance outcomes and inform policies for intervening when queue lengths signal potentially worsening system performance.
@@ 1 Introduction
A simulation model is a computational representation of a real-world system designed to study its behavior under various settings or scenarios. Simulation models are used extensively in fields such as engineering, economics, healthcare, and environmental science to predict and analyze outcomes of different scenarios without the need to experiment in the real world, which can be costly, time-consuming, or impractical. Outputs of a simulation model typically correspond to key performance indicators (KPIs) of interest to the decision maker, e.g., profit, throughput, or service level. For stochastic simulation models, simulating a given scenario generates outputs that vary from replication to replication, thus each scenario has an associated probability distribution describing the stochastic behavior of its outputs. This distribution can be approximated by the empirical distribution produced by the output data. When conducting simulation experiments, the user decides which scenarios are simulated and how many replications are run. Through the control of pseudo-random numbers, simulation experiments can be easily designed to produce output data that satisfies the standard assumptions of being independent and identically distributed, in contrast to most data obtained from the real world. Common tools for analyzing simulation output data include basic summary statistics (e.g., sample means, variance, and covariances) and visualization tools (e.g., histograms and boxplots). For problems with multiple KPIs, the multivariate empirical distribution produced by the data contains valuable information about system performance, but can be difficult to analyze and plot. To reveal important patterns and relationships that cannot be detected by conventional data analysis methods, we propose clustering the empirical distributions of simulated scenarios.
Clustering is an unsupervised learning approach that can help discover important patterns and relationships in complex datasets (James et al., 2013). In the context of simulation output data, clustering can identify scenarios with similar KPIs, or more precisely, similar output distributions. By understanding the characteristics of each cluster of distributions, a decision maker can draw meaningful comparisons of clusters and make informed decisions. We consider three important applications of clustering for enhancing simulation output analysis: anomaly detection, which involves identifying and investigating outliers in simulation outputs; pre-optimization, which involves formulating simulation-optimization problems and identifying promising initial solutions; and online monitoring, which involves tracking the system state over time and using classification methods to detect potentially undesirable system behavior and trigger appropriate actions or alerts. Our approach enhances decision making by extracting insights from complex simulation data with broad industry applications. For example, in call center optimization, our clustering algorithm can quickly distinguish staffing configurations that improve service levels and reduce costs. In healthcare operations, clustering can identify efficient resource allocation strategies that improve patient care and aid in evaluating trade-offs among metrics like wait times, resource utilization, and costs. In portfolio management, clustering can help with risk assessment and strategic planning by highlighting patterns in investment returns and enabling targeted interventions to prevent loss across different market conditions.
@@
We propose an agglomerative clustering method that uses the complete-linkage criterion for forming clusters. We choose agglomerative clustering for its flexibility, as it does not require specifying a predetermined number of clusters, and opt for complete linkage because it maintains compact and well-separated clusters. For the algorithm’s measure of dissimilarity between distributions, we choose the regularized Wasserstein distance. The (unregularized) Wasserstein distance quantifies the distance between two discrete probability distributions by the minimum amount of “work”—defined as the product of the probability mass that needs to be moved and the distance it needs to be transported—required to transform one distribution into the other (Villani, 2009). The Wasserstein distance is chosen over other metrics like Kullback-Leibler divergence or Jensen-Shannon divergence due to its ability to handle distributions with non-overlapping supports, as arises when working with continuous-valued empirical distributions. Another advantage of the Wasserstein distance is the notion of a barycenter, which acts like an average among distributions. Barycenters are particularly useful for post-hoc analysis, when desiring to summarize each cluster with a representative distribution (Peyré and Cuturi, 2019; Arjovsky et al., 2017; Cuturi, 2013). However, computing the Wasserstein distance between two discrete distributions entails solving a linear program (Villani, 2009), which could become computationally intensive when working with large datasets. The regularized Wasserstein distance, also known as the Sinkhorn distance, adds an entropic regularization term that promotes smoother transport plans, is easier to compute, and more stable, making it well suited for data-intensive applications (Cuturi, 2013). Our algorithm determines the optimal number of clusters based on the silhouette index (Shahapure and Nicholas, 2020), a centroid-free metric that evaluates the quality of a clustering in terms of its intra-cluster and inter-cluster distances.
@@
Clustering has been used for many purposes in many fields, such as analyzing gene expression in bioinformatics (Eisen et al., 1998), market segmentation in economics (Wedel and Kamakura, 2000), and textures and shapes in image processing (Pappas and Jayant, 1989). Hierarchical clustering techniques, specifically agglomerative clustering, build nested clusterings by iteratively merging the two most similar clusters (Murtagh and Contreras, 2012) and produce a dendrogram that depicts the nested clusterings at different levels (Jain et al., 1999). Adaptations like single-linkage, complete-linkage, and average-linkage offer great flexibility in how agglomerative clustering algorithms merge clusters (Jain, 2010). More recently, clustering has been employed in the study of stochastic simulation models, including for simulation optimization (Li et al., 2024; Zhang and Peng, 2024; Peng et al., 2018) and reducing model-form uncertainty (Abdallah et al., 2020). To the best of our knowledge, we are the first to propose agglomerative clustering of simulation output distributions and investigate important use cases.
@@
The Wasserstein distance is used extensively in machine learning applications, such as in computer vision and pattern recognition to robustly compare visual feature distributions (Rubner et al., 2000). For large-scale problems, variations like the regularized Wasserstein distance improve computational efficiency. For example, Benamou et al. (2015) employ iterative Bregman projections to efficiently solve regularized transportation problems, thereby improving scalability and accuracy. Our clustering framework uses the algorithm of Benamou et al. (2015) to calculate the regularized Wasserstein distance and regularized Wasserstein barycenters. For a more detailed overview of computational techniques for optimal transport and its practical applications, we direct the reader to Peyré and Cuturi (2019).
The Wasserstein distance has been applied in $k$-means clustering, particularly in the identification of market regimes (Horvath et al., 2021) and the analysis of financial data (Riess et al., 2023). Instead of relying on standard $k$-means, Henderson et al. (2015) proposed an extension to cluster one-dimensional, continuous-valued empirical distributions. However, as sample sizes grow, their EP-MEANS algorithm becomes less efficient and struggles to generalize to multivariate cases. A different approach was introduced by Irpino et al. (2014), who developed adaptive algorithms using squared Wasserstein distances, leading to improved performance in both synthetic and real-world scenarios. Meanwhile, Del Barrio et al. (2019) focused on enhancing robustness and efficiency by applying trimmed $k$-means on distributions with the squared Wasserstein distance. Zhuang et al. (2022) expanded $k$-means to the Wasserstein space, though they encountered issues with irregularity and non-robustness in barycenter-based clustering methods. Our hierarchical clustering algorithm addresses many of these limitations, particularly in terms of regularity and scalability. Another hierarchical clustering method based on optimal transport distance measures is offered by Chakraborty et al. (2020), though their work does not directly target the clustering of distributions as ours does.
@@ 2 Use Cases
Before presenting the algorithm itself, we first discuss the utility and versatility of clustering simulation output distributions through several use cases.
@@
In the context of simulation experiments, anomalies can be categorized as artificial or systemic. An artificial anomaly is typically associated with logic or coding errors within the simulation model, whereas a systemic anomaly is related to inherent features of the system. When using hierarchical clustering algorithms, anomalous output distributions can be identified by examining the dendrogram, the distances between clusters, or the cluster sizes (Loureiro et al., 2004). After identifying an anomalous output distribution, the decision maker would first scrutinize the simulation code to determine if it is an artificial anomaly. If the anomaly is not artificial, then he or she might investigate further by, for example, examining the marginal distributions, correlation matrices, and corresponding input variables. Hierarchical clustering algorithms such as ours are expected to be more stable for identifying outliers than non-hierarchical clustering methods whose clusterings strongly depend on the initialization of the clusters.
@@
In many practical situations, there are tradeoffs between KPIs, and the decision maker may be unable to articulate a priori what constitutes “good” versus “bad” system performance. By clustering output distributions and obtaining the barycenters, the decision maker can be presented with a more manageable number of distributions to compare. The decision maker could then conduct a series of pairwise comparisons of barycenters and thereby eliminate some clusters based on unformalized notions of preferred performance until a small number of clusters (or scenarios) remain. The clustering analysis can also help the decision maker specify which metrics should be modeled as objectives in a subsequent simulation-optimization problem and which should be treated as constraints. Achievable thresholds for the constraints can be set based on the observed performance outcomes of the simulated scenarios. Additionally, by examining the inputs associated with scenarios in a promising cluster, the decision maker can identify promising regions of the input space from which to initiate an optimization search, potentially leading to more rapid progress toward the optimal solution.
@@
This application concerns how the output of a simulation model is influenced by state variables, namely, those that evolve over time and can be observed, but not directly controlled, by the decision maker. We envisage an online monitoring framework in which clustering is performed offline and state variables are later tracked in real time and classification algorithms are utilized to help the decision maker anticipate changes in system performance. This approach involves a preliminary simulation experiment in which the scenarios correspond to different initial states, followed by the clustering of the generated outputs. When monitoring the system’s state online, classification algorithms can be used to predict the cluster to which an observed state’s output distribution may belong. Conversely, if the classification algorithm struggles to assign a state to a single cluster, such as a tie when using $k$-nearest neighbors, it could indicate that the system’s performance may change soon, potentially prompting intervention.
@@ 3 Clustering Simulation Output Distributions
Suppose there are $N$ scenarios under consideration, and for each Scenario $i$, $i=1,2,\ldots,N$, we obtain $n_{i}$ independent simulation replications. We make no assumptions about whether the replications performed at different systems are independent. In the numerical experiments, we examine the use of common random numbers (CRN) to induce positive correlation in the outputs from different systems and observe its effect on our clustering technique. Let $\mathbf{y}_{il}\in\mathbb{R}^{d}$ denote the vector output of the $l$th simulation replication at Scenario $i$ and let $\mu_{i}:=n_{i}^{-1}\sum_{l=1}^{n_{i}}\delta_{\mathbf{y}_{il}}$ denote the corresponding empirical distribution, i.e., a discrete probability distribution with support $\mathcal{Y}_{i}=\{\mathbf{y}_{i1},\ldots,\mathbf{y}_{in_{i}}\}$, where we ignore any duplicate values in the definition of $\mathcal{Y}_{i}$, and $\delta_{\mathbf{y}_{il}}$ is the Dirac delta function at $\mathbf{y}_{il}$.
We are interested in clustering the empirical distributions $\mu_{1},\mu_{2},\ldots,\mu_{N}$. Since there is no specific ordering among the distributions, we henceforth drop the subscript $i$ and denote the probability mass vector, support, and cardinality of the support of a generic empirical distribution $\mu$ as $\mathbf{p}_{\mu}$, $\mathcal{Y}_{\mu}$, and $M_{\mu}=|\mathcal{Y}_{\mu}|$, respectively.
@@ 3.1 Wasserstein Distance
Let $\Delta_{M}:=\left\{\mathbf{p}\in\mathbb{R}_{+}^{{M}}\colon\sum_{l=1}^{{M}}p_{l}=1\right\}$ denote the set of all possible probability mass vectors on a support of size $M$, where $\mathbb{R}_{+}^{{M}}$ denotes the set of all real-valued length-$M$ vectors having only non-negative components. For two empirical distributions $\mu$ and $\mu^{\prime}$ having probability mass vectors $\mathbf{p}_{\mu}\in\Delta_{M_{\mu}}$ and $\mathbf{p}_{\mu^{\prime}}\in\Delta_{M_{\mu^{\prime}}}$, respectively, the polytope of couplings is defined as
@@
The regularized Wasserstein distance (Genevay et al., 2018), on the other hand, can be computed in near-linear time (Cuturi, 2013) and is defined as
$W_{\lambda}(\mu,\mu^{\prime}):=\langle\mathbf{D},\boldsymbol{\boldsymbol{\gamma}_{\lambda}^{*}}\rangle,\quad\text{where}\quad\boldsymbol{\gamma}_{\lambda}^{*}:=\underset{\boldsymbol{\gamma}_{\lambda}\in\Pi(p_{\mu},p_{\mu^{\prime}})}{\operatorname{argmin}}\langle\mathbf{D},\boldsymbol{\gamma}_{\lambda}\rangle-\lambda E(\boldsymbol{\gamma}_{\lambda})$ (2)
where $\lambda>0$ is a regularization parameter, and $E(\boldsymbol{\gamma}_{\lambda})$ is the entropy of the transportation plan matrix $\boldsymbol{\gamma}_{\lambda}$, defined as $E(\boldsymbol{\gamma}):=-\sum_{l=1}^{M_{\mu}}\sum_{l^{\prime}=1}^{M_{\mu^{\prime}}}\boldsymbol{\gamma}_{ll^{\prime}}\log\boldsymbol{\gamma}_{ll^{\prime}}$, where we set $\boldsymbol{\gamma}_{ll^{\prime}}\log\boldsymbol{\gamma}_{ll^{\prime}}=0$ if $\boldsymbol{\gamma}_{ll^{\prime}}=0$. As $\lambda$ approaches 00, the optimal transportation plan matrix for (2), $\boldsymbol{\gamma}_{\lambda}^{*}$, becomes more sparse and approaches $\boldsymbol{\gamma}^{*}$ (Benamou et al., 2015). The entropic regularization term incentivizes $\boldsymbol{\gamma}_{\lambda}^{*}$ to be more diffuse than $\boldsymbol{\gamma}^{*}$, and this induced non-sparsity helps to stabilize the computation of $\boldsymbol{\gamma}_{\lambda}^{*}$ because (2) is a strongly convex program with a unique solution. An advantage of the regularized Wasserstein distance is that $\boldsymbol{\gamma}^{*}_{\lambda}$ can be calculated through an efficient iterative procedure involving matrix multiplications, as described in Algorithm 1.
Algorithm 1 Regularized Wasserstein distance (Benamou et al., 2015)
@@
2:Construct a matrix $\mathbf{Q}\in\mathbb{R}^{M_{\mu}\times M_{\mu^{\prime}}}$ having entries $Q_{ll^{\prime}}=e^{-\frac{\mathbf{D}_{ll^{\prime}}}{\lambda}}$ for $l=1,2,\ldots,M_{\mu}$ and $l^{\prime}=1,2,\ldots,M_{\mu^{\prime}}$.
@@
The stopping criteria in Algorithm 1 helps to control the computational cost and could involve setting a maximum number of iterations or stopping when the percentage change in the regularized Wasserstein distance is less than some threshold. The regularized Wasserstein distance plays a central role in our proposed clustering algorithm. Figure 1 demonstrates that the time required to calculate the Wasserstein distance increases exponentially in the support size, while the cost of computing the regularized Wasserstein distance increases nearly linearly.
Figure 1: Computational cost of calculating the Wasserstein distance and regularized Wasserstein distance between two samples generated with the same support size, varying between 1 and 3000, with calculations repeated 10 times for each support size. The variability was negligible compared to the difference in computational costs between the two distance measures.